1000 Litre milk contains some percent milk and some percent water. Now add some 70 Litre water to the above mixture. What is the percentage of water in the mixture?
Answer by stanbon(57290)

1000 Litre milk contains some percent milk and some percent water. Now add some 70 Litre water to the above mixture. What is the percentage of water in the mixture?
---
Assume the 1000 litres is x% milk.
Then it is (100-x)% water
Total water is [(100-x)/100]*1000 = 10(100-x) = 1000-10x litres
--------------------------------
Total liquid after adding 70 litres of water is 1070 litres.
-----
% of water in the final amount = (1000-10x)/1070 = (100-x)/107
===========================
Cheers,
Stan H.